When 4 litre container is filled at 27 degree Celsius the pressure inside is 850 millimetre of HG this container is placed in the ice bath what will be the pressure in it??

Grade:11

1 Answers

Vikas TU
4 years ago
Dear student 
Voume will be constant 
So, 
P1 / T1 = P2 /T2 
850 /300 = P2 /273 
P2 = 773 mm of Hg 
Hope this helps